OpenStudy (anonymous):

The population of a region is growing exponentially. There were 40 million people in 1990 (t = 0) and 70 million in 2000. Find an expression for the population at any time t, in years. Use the general exponential function and remember to use exact values. P(t)= (in millions) - What population would you predict for the year 2010? - What is the doubling time? (Round your answer to one decimal place.)

OpenStudy (anonymous):

Now to find the general exponential function, we need to find out some information. P(t)= Po x R^t Where: P= total population Po=initial population R= rate t= no of years after 1990

OpenStudy (anonymous):

We have the current info: P= 70 000 000 Po= 40 000 000 R=? t= 2000-1990=10years

OpenStudy (anonymous):

So sub the info into the equation to find the R.